# Diffusion-HMC: Parameter Inference with Diffusion Model driven Hamiltonian Monte Carlo

### v1

@@ IV.2.1 Robustness
To test this, we examine the extent to which our parameter estimates change relative to the predictions of the parameter inference network in [48]. The neural network in [48] is trained to predict the mean and the standard deviation of the parameters given an input dark matter field. In the leftmost column of Figure  7, we compare the sample predictions obtained from the diffusion model (green) relative to those obtained from the parameter inference network (red) on 5 fields ($x_{0}$). In the next two columns, we add increasing levels of $\mathcal{N}(0,\sigma)$ noise with $\sigma\in[0.01,0.02]$, and examine the estimates for both approaches. We perform this experiment in a noise-agnostic setting, i.e. we assume that we do not know the level of noise a priori and do not modify the parameter inference approach for the diffusion model. We use the same $T_{\rm{MAX}}=20$ and use 200 samples for the estimates.
@@ V Conclusion
We then directed our attention to the inverse problem of parameter inference and disentangled the contribution of each term $L_{t}$ in the expression for the Variational Lower Bound to constraints on cosmological parameters. Our findings reveal that the strength of the constraints decrease as $t$ increases. [21] explored this question in the context of using the diffusion model’s conditional VLB terms for classification and found that intermediate timesteps had the highest accuracy when only a single timestep is used. The difference in which terms contain the most information about the conditioning attribute (cosmology / class) is interesting, and could be partly explained by the difference in their formulation and weighting of the VLB terms and in how changes in the conditioning attribute affect an image in the two settings. Modulating cosmology modifies global attributes of the fields, such as their intensity and power spectra, as seen in Figure  2, while the information that distinguishes different breeds of pets from each other tends to be relatively localized in the bounding box containing the animal.
@@
Lastly, we demonstrate that the diffusion model likelihood confers the Diffusion-HMC constraints with greater robustness to the addition of noise to the input image relative to the behavior of a parameter inference network. This echoes the behavior of diffusion models in other discriminative tasks such as [21, 5, 31], where diffusion model based classifiers have been shown to possess higher robustness to adversarial examples or perturbations. This is a pertinent finding since [17] showed that the powerful constraints derived from neural network based parameter inference may often come at the cost of their susceptibility to slight perturbations that the canonical two-point correlation based analyses are impervious to.
@@ VII Acknowledgements
Appendix
Refer to captionRefer to caption
@@
$q(x_{t}|x_{t+1},x_{0})=\mathcal{N}(\tilde{\mu_{t}}(x_{t+1}x_{0}),\tilde{\beta}_{t})$
$\tilde{\mu_{t}}(x_{t+1}x_{0})=\frac{\sqrt{\bar{\alpha}_{t-1}\beta_{t}}}{1-\bar{\alpha}_{t}}x_{0}+\frac{\sqrt{\alpha_{t}(1-\bar{\alpha}_{t-1})}}{1-\bar{\alpha}_{t}}x_{t+1}\text{ and }\tilde{\beta}_{t}=\beta_{t}\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}$
@@
.2 Parameter Estimation Baselines
@@
.3 Summary Statistics
.3.1 Reduced Chi-squared Statistics

### v2

@@ IV.3 Robustness
To test this, we examine the extent to which our parameter estimates change relative to the predictions of the parameter inference network in [55]. The neural network in [55] is trained to predict the mean and the standard deviation of the parameters given an input dark matter field. In the leftmost column of Figure  7, we compare sample predictions obtained from the diffusion model (Diffusion-HMC) relative to those obtained from the parameter inference network (InfNet) on 5 fields ($x_{0}$). In the next two columns, we add increasing levels of $\mathcal{N}(0,\sigma)$ noise with $\sigma\in[0.01,0.02]$, and examine the estimates for both approaches. We perform this experiment in a noise-agnostic setting, i.e. we assume that we do not know the level of noise a priori and do not modify either the Diffusion-HMC approach or the parameter inference network based approach. We use the same $T_{\rm{MAX}}=20$ and use 200 samples for the Diffusion-HMC estimates for all fields. Although the noise perturbations to the field are visually imperceptible, and correspond to a change of less than $\sim 4-6\%$ to the power spectrum at the smallest scales, the predictions of the parameter inference network are significantly ($\sim 6-23\%$) disturbed. Indeed, we find that the Diffusion-HMC based constraints are perturbed far less than the InfNet constraints. The Diffusion-HMC $\sigma_{8}$ constraints are also more robust than its $\Omega_{m}$ constraints.
@@ V Conclusion
We then directed our attention to the inverse problem of parameter inference and disentangled the contribution of each term $L_{t}$ in the expression for the Variational Lower Bound to constraints on cosmological parameters. Our findings reveal that the strength of the constraints decrease as $t$ increases. [6] and [24] explored this question in the context of using the diffusion model’s conditional VLB terms for classification. [6] also found a negative exponential weighting of the terms to be optimum while [24] found that intermediate timesteps had the highest accuracy when only a single timestep is used. The difference in which terms contain the most information about the conditioning attribute (cosmology / class) is interesting, and could be partly attributed to the difference in their formulation and weighting of the VLB terms and in how changes in the conditioning vector affect an image in different settings. Modulating cosmology modifies global attributes of the fields, such as their intensity and power spectra, as seen in Figure  2, while the information that distinguishes different breeds of pets from each other tends to be relatively localized in the bounding box containing the animal.
@@
Lastly, we demonstrate that the diffusion model likelihood confers the Diffusion-HMC constraints with greater robustness against the addition of small amounts of noise to the input image relative to the behavior of a parameter inference network. This echoes the behavior of diffusion models in other discriminative tasks such as [24, 5, 36], where diffusion model based classifiers have been shown to possess higher robustness to adversarial examples or perturbations. This is a pertinent finding since [18] showed that the powerful constraints derived from neural network based parameter inference may often come at the cost of their susceptibility to slight perturbations that the canonical two-point correlation based analyses are impervious to. We find that our diffusion-model based parameter inference approach enables more noise-robust field-level inference. It would be interesting to further explore the differences in inductive biases learned by discriminative networks (e.g.: [55, 43]) relative to generative models repurposed for discriminative tasks.
@@ VII Acknowledgements
Appendix A Parameter Inference
@@
$q(x_{t}|x_{t+1},x_{0})=\mathcal{N}(\tilde{\mu_{t}}(x_{t+1},x_{0}),\tilde{\beta}_{t})$
$\tilde{\mu_{t}}(x_{t+1},x_{0})=\frac{\sqrt{\bar{\alpha}_{t-1}\beta_{t}}}{1-\bar{\alpha}_{t}}x_{0}+\frac{\sqrt{\alpha_{t}(1-\bar{\alpha}_{t-1})}}{1-\bar{\alpha}_{t}}x_{t+1}\text{ and }\tilde{\beta}_{t}=\beta_{t}\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}$
@@ A.1 HMC Convergence
A.2 Parameter Estimation Baselines
@@
A.3 Additional Robustness Tests
Dropping the Prior:
In Figure 10, we explore robustness without the prior in the HMC setting, for the noise levels in Figure 7 as well as with the addition of more noise. For small amounts of noise, ($\sigma$ = 0.01, 0.02) removing the prior doesn’t affect the bias of the Diffusion-HMC-inferred parameters since the numbers in Figure 10 are identical / close to those in Figure 7. For noise with a scale of 0.05, the Diffusion-HMC constraints are more perturbed for $\Omega_{m}$ but more robust for $\sigma_{8}$ relative to the neural network baseline. A standard deviation of 0.05 is equivalent to the diffusion noising timesteps between the 1st and the 2nd timestep, while {0.01, 0.02} are less than the first timestep. Note, we plot the power spectra of the ‘true’ (linear) mass density field in these figures, to be consistent with the power spectra in Figures 1 and 2. However, the noise is added to the log of the field. Thus while the effect on the linear power spectrum is mild (around 5% at the smallest scales) for $\sigma=0.05$ the effect on the power spectrum of the log field is around 35% at the smallest scales.
Dropping Initial Timesteps:
In the left panel of Figure 5, we observed that the isolated effect of the smallest timesteps also had the strongest constraints. We thus ask the question, if we perform parameter inference in a setting where there is some knowledge of the amount of noise added, can dropping timesteps confer additional robustness? In Figure 11, we add a comparison for the noising amounts corresponding to 0.02 and 0.05 and drop the first 2 timesteps. The rest of the setting is the same as in Figure 10 i.e. we also drop the prior. We find that dropping these timesteps confers greater robustness on the estimates, without noticeably reducing confidence. While we defer a more rigorous investigation to future work, the prospect of using the knowledge of the amount of noise added to strategically drop timesteps could be of interest, reminiscent of scale-dependent analyses and scale cuts in other cosmological analysis methods [10, 22].
Figure 10: Robustness comparison without a prior applied in the HMC case.
Figure 11: Robustness comparison with the first 2 noising timesteps dropped and without a prior applied in the HMC case.
Appendix B Summary Statistics
B.1 Reduced Chi-squared Statistics